At T= 100 K, 1.00 mol of CO2 occupies a volume of 1.00 L. Calculate the pressures given by the ideal gas equation and the van der Waals equation. (a= 3.610 atm dm6 mol-2) and (b = 4.29 dm3 mol-1).


1
Expert's answer
2021-05-17T04:19:43-0400

For Ideal gas .

PV = nRT .

P= nRT / V .

P= 1 × 0.0821 × 100 / 1 .

P= 0.000821 atm .



For Real gas .

( P + an^2 /v^2 ) ( v- nb ) = nR T .


( P + 3.610 × 1 ) ( 1 - 4.29 ) = 1 × 0.0821 × 100


Then find p .